# Diffusion posterior sampling for simulation-based inference in tall data settings

### 11 Apr 2024 v1

@@ 3.1 Exact computation of the tall data posterior score
Let $x^{\star}_{1:n}=\{x^{\star}_{1},\dots,x^{\star}_{n}\}\sim_{\rm{iid}}p(x\mid\theta^{\star})$ be observations sampled using the same parameter $\theta^{\star}\in\mathbb{R}^{m}$. Our goal is to sample from the tall data posterior $p(\theta\mid x^{\star}_{1:n})$ via the backward sampling Markov chain defined in (5), while only relying on a score estimate $s_{\phi}(\theta,x,\alpha_{t})$ of $\nabla_{\theta}\log p_{t}(\theta\mid x)$ of the diffused posterior for a single observation. To do so, we need to build an approximation of the diffused tall data posterior score
$\nabla_{\theta_{t}}\log p_{t}(\theta_{t}\mid x^{\star}_{1:n})=\nabla_{\theta_{t}}\log\int p(\theta\mid x^{\star}_{1:n})q_{t|0}(\theta_{t}|\theta)\mathrm{d}\theta$
that solely depends on $\nabla_{\theta}\log p_{t}(\theta\mid x^{\star}_{j})$, for all $j\in[1,n]$. Using (1) we can write
@@
Given the diffused prior $p^{\lambda}_{t}(\theta)=\int\lambda(\theta_{0})q_{t|0}(\theta|\theta_{0})\mathrm{d}\theta_{0}$ [^1] The diffused prior can be computed analytically for simple (e.g. uniform or Gaussian) prior distributions as in Sharrock et al., (2022)., we now introduce the following backward transition kernels obtained via Bayes’ rule:
@@
Rearranging terms in equation (7) and replacing them with the quantities in (8), gives us
@@
$=\int\left(p^{\lambda}_{0|t}(\theta_{0}|\theta)p^{\lambda}_{t}(\theta)\right)^{1-n}\prod_{j=1}^{n}p_{0|t}(\theta_{0}|\theta,x^{\star}_{j})p_{t}(\theta\mid x^{\star}_{j})\mathrm{d}\theta_{0}$
$=\left(p^{\lambda}_{t}(\theta)^{1-n}\prod_{j=1}^{n}p_{t}(\theta\mid x^{\star}_{j})\right)L_{\lambda}(\theta,x^{\star}_{1:n})\,,$
with $L_{\lambda}(\theta,x^{\star}_{1:n})\vcentcolon=\int p^{\lambda}_{0|t}(\theta_{0}|\theta)^{1-n}\prod_{j=1}^{n}p_{0|t}(\theta_{0}|\theta,x^{\star}_{j})\mathrm{d}\theta_{0}\,.$ The associated Fisher score therefore writes
@@
The first two terms in the above equation are known: the prior score can be computed analytically in most cases [^2] See Appendix D for the Gaussian and Uniform cases. If not, it can be learned via the classifier-free guidance approach (Ho and Salimans,, 2022), at the same time as the posterior score (more details and experimental results are provided in Appendix K). and the (single) posterior scores are approximated by evaluating the learned score model $s_{\phi}(\theta,x,\alpha_{t})$ in every $x^{\star}_{j}$. This leaves us with the last term, that we still need to approximate.
@@ 3.2 Second order approximation of $\log L_{\lambda}$
In this section, we consider a Gaussian approximation of the backward kernels defined in (8) via
$\hat{p}_{0|t}(\theta_{0}|\theta,x)$ $=\mathcal{N}(\theta_{0};\mathrm{\mu}_{t}(\theta,x),\Sigma_{t}(\theta,x))$ (9)
$\text{and}\quad\hat{p}^{\lambda}_{0|t}(\theta_{0}|\theta)$ $=\mathcal{N}(\theta_{0};\mathrm{\mu}_{t,\lambda}(\theta),\Sigma_{t,\lambda}(\theta))\,,$ (10)
where $\mathrm{\mu}_{t}(\theta,x),\ \mathrm{\mu}_{t,\lambda}(\theta)$ and $\Sigma_{t}(\theta,x),\ \Sigma_{t,\lambda}(\theta)$ are the means and covariance matrices of the backward processes respectively associated with the diffused posterior $p_{t}(\theta\mid x)$ and prior $p_{t}^{\lambda}(\theta)$ distributions. [^3] Note that ${\mathrm{\mu}_{t}(\theta)=\mathbb{E}\left[\theta_{0}\mid\theta\right]}$ and ${\Sigma_{t}(\theta)=\mathbb{E}\left[\left(\theta_{0}-\mathrm{\mu}_{t}(\theta)\right)\left(\theta_{0}-\mathrm{\mu}_{t}(\theta)\right)^{\top}\mid\theta\right]}$. They are explicitly known for certain distribution choices and are respectively functions of the score and its derivatives; see Boys et al., (2023). This leads us to the following estimator of $\log L_{\lambda}(\theta,x^{\star}_{1:n})$:
$\ell_{\lambda}(\theta,x^{\star}_{1:n})=\log\int\hat{p}^{\lambda}_{0|t}(\theta_{0}|\theta)^{1-n}\prod_{j=1}^{n}\hat{p}_{0|t}(\theta_{0}|\theta,x_{j})\mathrm{d}\theta_{0}\,.$ (11)
@@
For all $\theta\in\mathbb{R}^{m}$, let $\Lambda(\theta)=\sum_{j=1}^{n}\Sigma^{-1}_{t,j}(\theta)+(1-n)\Sigma^{-1}_{t,\lambda}(\theta)$. Assume that $\Lambda(\theta)$ is positive definite. Then, for all $\theta\in\mathbb{R}^{m}$ and $x^{\star}_{i}\in\mathbb{R}^{d}$, $1\leq i\leq n$,
@@
where
$\zeta(\mu,\Sigma)=-\left(m\log 2\pi-\log|\Sigma^{-1}|+\mu^{\top}\Sigma^{-1}\mu\right)/2\,,$
and $\zeta_{j}(\theta)=\zeta(\mathrm{\mu}_{t,j}(\theta),\Sigma_{t,j}(\theta))$ for all $j\in[1:n]$, $\zeta_{\lambda}(\theta)=\zeta(\mathrm{\mu}_{t,\lambda}(\theta),\Sigma_{t,\lambda}(\theta))$ and $\zeta_{\operatorname{all}}(\theta)=\zeta(\Lambda(\theta)^{-1}\eta(\theta),\Lambda(\theta)^{-1})$ with $\eta(\theta)=\sum_{j=1}^{n}\Sigma^{-1}_{t,j}(\theta)\mathrm{\mu}_{t,j}(\theta)+(1-n)\Sigma^{-1}_{t,\lambda}(\theta)\mathrm{\mu}_{t,\lambda}(\theta)$.
@@
For all $\theta\in\mathbb{R}^{m}$ and $x^{\star}_{i}\in\mathbb{R}^{d}$, $1\leq i\leq n$, the full gradient can be written as
$\nabla_{\theta}\log p_{t}(\theta\mid x^{\star}_{1:n})=\Lambda(\theta)^{-1}\sum_{j=1}^{n}\Sigma^{-1}_{t,j}(\theta)\nabla_{\theta}\log p_{t}(\theta\mid x_{j})$
$\quad+(1-n)\Lambda(\theta)^{-1}\Sigma^{-1}_{t,\lambda}(\theta)\nabla_{\theta}\log p^{\lambda}_{t}(\theta){}+F(\theta,x^{\star}_{1:n})\,,$
where $F(\theta,x_{1:n})=0$ if $\nabla_{\theta}\Sigma_{t,j}(\theta)=0$ for all $j\in[1:n]$ and $\nabla_{\theta}\Sigma_{\lambda,t}(\theta)=0$.
We now focus on the particular choice of $\Sigma_{t,i}(\theta)$ that we use in our approximation. It can be shown that $\Sigma_{t,i}(\theta_{t})=\nabla_{\theta_{t}}\mu_{t,i}(\theta_{t})$ (Boys et al.,, 2023). This corresponds to Algorithm 2, which we refer to as JAC. However, this approach possesses two main drawbacks. The first is that it scales poorly with the dimension; indeed, we need to calculate a $m\times m$ matrix which is prohibitive for large $m$. The second is that we have to take the derivative w.r.t. the inputs of the score neural network, which is known to be unstable. Note that as proposed in Boys et al., (2023), we do not propagate gradients through $\Sigma_{t,i}(\theta_{t})$, rendering $F(\theta_{t},x_{1:n})=0$ in Lemma 3.2.
As an alternative, we consider a constant approximation of the covariance matrix. To do so, we start by noting that in the case where $q_{\operatorname{data}}$ is a Gaussian distribution with mean $\mu_{0}$ and variance $\Sigma_{0}$, we have $\Sigma_{t}=(\Sigma_{0}^{-1}+\alpha_{t}(1-\alpha_{t})^{-1}\mathbf{I}_{m})^{-1}$, see Appendix D. Note that choosing $\Sigma_{0}=\mathbf{I}_{m}$ results in the approximation proposed by Song et al., (2023). We use the formula for the Gaussian case as an approximation of the real covariance matrix, yielding Algorithm 1 which we refer to as GAUSS. To do so, we need to first estimate $\Sigma_{0}$ for each $x$, which we do by first running a DDIM with a small number of iterations ($100$) for each $i$ and using the empirical covariance matrix of the samples as $\Sigma_{0}$. Note that in this case, we also have $F(\theta_{t},x_{1:n})=0$ in Lemma 3.2.
Algorithm 1 GAUSS (Gaussian approximation)
@@
Algorithm 2 JAC (Jacobian approximation)
@@ 4 Experiments
In this section, we investigate the performance of GAUSS and JAC for different tasks from SBI literature with increasing difficulty. We take F-NPSE (Geffner et al.,, 2023) as a baseline for comparisons, which uses unadjusted Langevin dynamics (ULD) with $L=5$ steps and $\tau=0.5$. We have decided to not compare our methods to augmented versions of NPE, NLE, and NRE because this has already been done in Sharrock et al., (2022) and Geffner et al., (2023) with the score-diffusion framework demonstrating clearly superior performance. Of course, this choice reduces the range of comparisons of our experimental section, but allows us to focus solely on the best alternative method from current literature. For GAUSS and JAC we sample with the backward Markov chain defined in Section 2 with $\sigma^{2}_{t}=\eta^{2}(1-\alpha_{t-1})/(1-\alpha_{t})(1-\alpha_{t}/\alpha_{t-1})$ where $\eta=0.2,0.5,0.8,1$ for a number of steps $T=50,150,400,1000$ respectively. We use a uniform scheduling $\{t_{i}=i/T\}_{i=1}^{T}$. The code reproducing all experiments is available in the following repository: https://github.com/JuliaLinhart/diffusions-for-sbi.
@@ 4.1 Gaussian toy models
We consider two toy examples for which the analytic form of the posterior and corresponding score distributions are known. Our first example is a multivariate Gaussian simulator defined by $p(x\mid\theta)=$ $\mathcal{N}(x;\theta,(1-\rho)\mathbf{I}_{m}+\rho\mathbf{1}_{m})$ with correlation factor $\rho=0.8$. The second is a Mixture of Gaussians (GMM) $p(x\mid\theta)=0.5~{}\mathcal{N}(x;\theta,2.25\Sigma)+0.5~{}\mathcal{N}(x;\theta,\Sigma/9)$, where $\Sigma$ is a diagonal matrix with values increasing linearly between $0.6$ and $1.4$, following  Geffner et al., (2023). Both examples are carried out with a Gaussian prior $\lambda(\theta)=\mathcal{N}(\theta;0,\mathbf{I}_{m})$ in dimension $m=10$. Here, the samples of the true posterior for multiple observations are obtained either directly for Gaussian or via Metropolis Adjusted Langevin (MALA) for GMM (see Appendix E for further details). We use the sliced Wasserstein (sW) distance to measure how close the obtained posterior samples are to the true samples. [^4] In order to account for the finite-sample effects on sW, we considered a normalized version of the distance by removing the expected sW over different sample sets from the true distribution. The sW is computed with $10^{4}$ slices, on $10^{3}$ samples and results are reported over 5 different seeds.
Consider the following approximate score model
$\tilde{\mathrm{s}}_{\psi}(\theta,x,\alpha_{t})=\mathrm{s}_{\psi}(\theta,x,\alpha_{t})+\epsilon(1-\alpha_{t})r(\theta,x,\alpha_{t})~{},$
where $\epsilon\geq 0$, $\mathrm{s}_{\psi}(\theta,x,\alpha_{t})$ is the analytical posterior score and $r$ is a randomly initialized neural net with outputs in range $[-1,1]$, see Appendix D. This construction leads to an error of $\epsilon$ over the “noise predictor” network defined as $-\tilde{\mathrm{s}}_{\psi}(\theta,x,\alpha_{t})/(1-\alpha_{t})$, which is the neural network that one actually optimizes when training a score model. Note that unlike the Gaussian case, Tweedie’s approximation is not exact for GMM, i.e. $p_{0|t|}(\theta_{0}|\theta_{t})$ is not a Gaussian density for all $\theta_{t}$ and all $t\in[1:T]$. Therefore, this example allows us to quantify the effect of Tweedie’s approximation while still controlling the score estimation error (through $\epsilon$).
@@
Table 1 displays the total running time and sW for each algorithm on the Gaussian example (Table 3 in Appendix G shows similar results for GMM). Based on this table, we consider “equivalent time settings”, namely we run our algorithm with $400$ and $1000$ steps for JAC and GAUSS respectively and LANGEVIN for $400$ steps. We can see that our algorithm yields smaller sW than the equivalent Langevin sampler while accounting for $5$ times less neural network evaluation (NNE), thus $5$ times faster. We show in Appendix G the differences in speed for each tested setting.
@@
Figure 2: Normalized sliced Wasserstein (sW) distance as a function of the number of observations $n$ for the Gaussian and GMM toy examples, for each algorithm (GAUSS, JAC and LANGEVIN) and with different levels of $\epsilon$. Mean and std over 5 different seeds.

### 07 Jun 2024 v2

@@ 3 Diffusion posterior sampling for tall data
In this section, we propose a new algorithm that approximately samples from the backward diffusion process associated with the factorized tall data posterior, thereby eliminating the need of costly and unstable Langevin steps, while using only individual scores obtained for each observation.
@@ 3.1 Exact computation of the tall data posterior score
Let $x^{\star}_{1:n}=\{x^{\star}_{1},\dots,x^{\star}_{n}\}\sim_{\rm{iid}}p(x\mid\theta^{\star})$ be observations sampled using the same parameter $\theta^{\star}\in\mathbb{R}^{m}$. Our goal is to sample from the tall data posterior $p(\theta\mid x^{\star}_{1:n})$ via the backward sampling Markov chain defined in (5), while only relying on a score estimate $s_{\phi}(\theta,x,\alpha_{t})\approx\nabla_{\theta}\log p_{t}(\theta\mid x)$ of the diffused posterior for a single observation $x$. To do so, we need to build an approximation of the diffused tall data posterior score $\nabla_{\theta_{t}}\log p_{t}(\theta_{t}\mid x^{\star}_{1:n})=\nabla_{\theta_{t}}\log\int p(\theta\mid x^{\star}_{1:n})q_{t|0}(\theta_{t}|\theta)\mathrm{d}\theta$ that solely depends on $\nabla_{\theta}\log p_{t}(\theta\mid x^{\star}_{j})$, for all $j\in[1,n]$. Using (1) we can write
@@
Given the diffused prior $p^{\lambda}_{t}(\theta)=\int\lambda(\theta_{0})q_{t|0}(\theta|\theta_{0})\mathrm{d}\theta_{0}$, we now introduce the following backward transition kernels obtained via Bayes’ rule:
@@
Rearranging terms in Equation (7) and using (8) yields
@@
$=L_{\lambda}(\theta,x^{\star}_{1:n})p^{\lambda}_{t}(\theta)^{1-n}\prod_{j=1}^{n}p_{t}(\theta\mid x^{\star}_{j})\,,$
with $L_{\lambda}(\theta,x^{\star}_{1:n}):=\int p^{\lambda}_{0|t}(\theta_{0}|\theta)^{1-n}\prod_{j=1}^{n}p_{0|t}(\theta_{0}|\theta,x^{\star}_{j})\mathrm{d}\theta_{0}$. The corresponding score writes
@@
The first two terms in the above equation are tractable: the prior score can be computed analytically in most cases [^1] See Appendix D for the Gaussian and Uniform cases. If not, it can be learned via the classifier-free guidance approach (Ho and Salimans,, 2021), at the same time as the posterior score (more details and experimental results are provided in Appendix L.1) and the (single) posterior scores are approximated by evaluating the learned score model $s_{\phi}(\theta,x,\alpha_{t})$ at every $x^{\star}_{j}$. This leaves the last term that integrates all backward kernels for each $x^{\star}_{j}$ which needs to be estimated. Note that this term is missing in the score formula from Equation (6) and is the reason why Langevin steps are required in F-NPSE.
@@ 3.2 Second order approximation of $\log L_{\lambda}$
We consider a Gaussian approximation of the backward kernels defined in (8) via
$\hat{p}_{0|t}(\theta_{0}|\theta,x)=\mathcal{N}(\theta_{0};\mathrm{\mu}_{t}(\theta,x),\Sigma_{t}(\theta,x))\quad\text{and}\quad\hat{p}^{\lambda}_{0|t}(\theta_{0}|\theta)=\mathcal{N}(\theta_{0};\mathrm{\mu}_{t,\lambda}(\theta),\Sigma_{t,\lambda}(\theta))\,,$ (9)
[…]
$\ell_{\lambda}(\theta,x^{\star}_{1:n})=\log\int\hat{p}^{\lambda}_{0|t}(\theta_{0}|\theta)^{1-n}\prod_{j=1}^{n}\hat{p}_{0|t}(\theta_{0}|\theta,x^{\star}_{j})\mathrm{d}\theta_{0}\,.$ (10)
@@
For all $\theta\in\mathbb{R}^{m}$, let $\Lambda(\theta)=\sum_{j=1}^{n}\Sigma^{-1}_{t,j}(\theta)+(1-n)\Sigma^{-1}_{t,\lambda}(\theta)$. Assume that $\Lambda(\theta)$ is positive definite. Then, for all $\theta\in\mathbb{R}^{m}$ and $x^{\star}_{j}\in\mathbb{R}^{d}$, $1\leq j\leq n$,
@@
where $\zeta(\mu,\Sigma)=-\left(m\log 2\pi-\log|\Sigma^{-1}|+\mu^{\top}\Sigma^{-1}\mu\right)/2$, and $\zeta_{j}(\theta)=\zeta(\mathrm{\mu}_{t,j}(\theta),\Sigma_{t,j}(\theta))$ for all $j\in[1:n]$, $\zeta_{\lambda}(\theta)=\zeta(\mathrm{\mu}_{t,\lambda}(\theta),\Sigma_{t,\lambda}(\theta))$ and $\zeta_{\operatorname{all}}(\theta)=\zeta(\Lambda(\theta)^{-1}\eta(\theta),\Lambda(\theta)^{-1})$ with $\eta(\theta)=\sum_{j=1}^{n}\Sigma^{-1}_{t,j}(\theta)\mathrm{\mu}_{t,j}(\theta)+(1-n)\Sigma^{-1}_{t,\lambda}(\theta)\mathrm{\mu}_{t,\lambda}(\theta)$.
@@
For all $\theta\in\mathbb{R}^{m}$ and $x^{\star}_{j}\in\mathbb{R}^{d}$, $1\leq j\leq n$, the full gradient can be written as
$\nabla_{\theta}\log p_{t}(\theta\mid x^{\star}_{1:n})=\Lambda(\theta)^{-1}\sum_{j=1}^{n}\Sigma^{-1}_{t,j}(\theta)\nabla_{\theta}\log p_{t}(\theta\mid x^{\star}_{j})\\ +(1-n)\Lambda(\theta)^{-1}\Sigma^{-1}_{t,\lambda}(\theta)\nabla_{\theta}\log p^{\lambda}_{t}(\theta){}+F(\theta,x^{\star}_{1:n})\,,$ (12)
where $F(\theta,x_{1:n})=0$ if $\nabla_{\theta}\Sigma_{t,j}(\theta)=0$ for all $1\leq j\leq n$ and $\nabla_{\theta}\Sigma_{\lambda,t}(\theta)=0$.
Formula (12) defines our approximation of the diffused tall posterior score and depends solely on the individual (prior and posterior) scores. We now focus on the computation of the backward covariance matrices $\Sigma_{t,\lambda}(\theta)$ and $\Sigma_{t,j}(\theta)$ with $1\leq j\leq n$, while keeping in mind that we want $F(\theta,x_{1:n})$ to be zero. First, note that for some prior choices, $\Sigma_{t,\lambda}(\theta)$ can be computed analytically; otherwise, we use the same strategy as for the intractable posterior case. This is represented by prior_fn in Algorithms 1 and 2.
From (Boys et al.,, 2023), it can be shown that $\Sigma_{t,j}(\theta)=\alpha_{t}^{-1/2}(1-\alpha_{t})\nabla_{\theta}\mu_{t,j}(\theta)$, where $\mu_{t,j}(\theta_{t})$ is completely defined with the posterior score as in Equation (4). $\Sigma_{t,j}(\theta_{t})$ can therefore be approximated by directly taking the Jacobian of the learned score function $s_{\phi}(\theta,x^{\star}_{j},\alpha_{t})$. As proposed in Boys et al., (2023), we do not propagate gradients through $\Sigma_{t,i}(\theta)$, rendering $F(\theta,x_{1:n})=0$ in Lemma 3.2. This corresponds to Algorithm 2, which we refer to as JAC. But this approach has two main drawbacks. First, we need to calculate a $m\times m$ matrix which is prohibitive for large $m$. Second, we have to take the derivative w.r.t. the inputs of the score neural network, which is known to be unstable.
As an alternative, we consider a constant approximation of the covariance matrix. To do so, we start by noting that in the case where $q_{\operatorname{data}}$ is a Gaussian distribution with mean $\mu_{0}$ and variance $\Sigma_{0}$, we have $\Sigma_{t}=(\Sigma_{0}^{-1}+\alpha_{t}(1-\alpha_{t})^{-1}\mathbf{I}_{m})^{-1}$, see Appendix D. Note that choosing $\Sigma_{0}=\mathbf{I}_{m}$ results in the approximation proposed by Song et al., (2023). We use the formula for the Gaussian case as an approximation of the real covariance matrix $\Sigma_{t,j}$, yielding Algorithm 1 which we refer to as GAUSS. To do so, we need to first estimate $\Sigma_{0,j}$ for each $x^{\star}_{j}$, which we do by first running a DDIM with a small number of iterations ($100$) for each $j$ and using the empirical covariance matrix of the samples as $\Sigma_{0,j}$. Note that in this case, we also have $F(\theta,x_{1:n})=0$ in Lemma 3.2.
Algorithm 1 GAUSS
(Gaussian approximation)
@@
Algorithm 2 JAC
(Jacobian approximation)
@@ 4 Experiments
We investigate the performance of GAUSS and JAC for different tasks from the SBI literature with increasing difficulty. We choose F-NPSE (Geffner et al.,, 2023) as a baseline for comparisons, which uses unadjusted Langevin dynamics (ULD) with $L=5$ steps and $\tau=0.5$. We have decided to not compare our methods to augmented versions of NPE, NLE, and NRE because it has already been shown that the score-diffusion framework demonstrates clear superior performance  (Sharrock et al.,, 2022; Geffner et al.,, 2023; Gloeckler et al.,, 2024). For GAUSS and JAC we sample with the backward Markov chain defined in Section 2 with $\sigma^{2}_{t}=\eta^{2}(1-\alpha_{t-1})/(1-\alpha_{t})(1-\alpha_{t}/\alpha_{t-1})$ where $\eta=0.2,0.5,0.8,1$ for a number of steps $T=50,150,400,1000$ respectively. We use a uniform scheduling $\{t_{i}=i/T\}_{i=1}^{T}$. To evaluate the accuracy of the sampling algorithms, we compute the sliced Wasserstein (sW) distance between estimated and true posterior samples. Further implementation details can be found in Appendix E and the code reproducing all experiments is available in the following repository: https://github.com/JuliaLinhart/diffusions-for-sbi.
@@ 4.1 Gaussian toy models
We consider two toy examples for which the analytic form of the posterior and corresponding score distributions are known. Our first example is a multivariate Gaussian simulator $p(x\mid\theta)=\mathcal{N}(x;\theta,(1-\rho)\mathbf{I}_{m}+\rho\mathbf{1}_{m})$ with correlation factor $\rho=0.8$. The second is a Mixture of Gaussians (GMM) $p(x\mid\theta)=0.5~{}\mathcal{N}(x;\theta,2.25\Sigma)+0.5~{}\mathcal{N}(x;\theta,\Sigma/9)$, where $\Sigma$ is a diagonal matrix with values increasing linearly between $0.6$ and $1.4$, following  Geffner et al., (2023). Both examples are carried out with a Gaussian prior $\lambda(\theta)=\mathcal{N}(\theta;0,\mathbf{I}_{m})$ in dimension $m=10$. Here, the samples of the true posterior with multiple observations can be obtained either directly for the Gaussian model or via Metropolis Adjusted Langevin (MALA) for GMM (see Appendix E for further details).
Consider the following approximate score model $\tilde{\mathrm{s}}_{\psi}(\theta,x,\alpha_{t})=\mathrm{s}_{\psi}(\theta,x,\alpha_{t})+\epsilon(1-\alpha_{t})r(\theta,x,\alpha_{t})~{},$ where $\epsilon\geq 0$, $\mathrm{s}_{\psi}(\theta,x,\alpha_{t})$ is the analytical posterior score and $r$ is a randomly initialized neural net with outputs in range $[-1,1]$, see Appendix D. This construction leads to an error of $\epsilon$ over the “noise predictor” network defined as $-\tilde{\mathrm{s}}_{\psi}(\theta,x,\alpha_{t})/(1-\alpha_{t})$, which is the neural network that one actually optimizes when training a score model. Note that unlike the Gaussian case, the Gaussian approximation from section 3.2 is not exact for GMM, i.e. $p_{0|t|}(\theta_{0}|\theta_{t})$ is not a Gaussian density for all $\theta_{t}$ and all $t\in[1:T]$. Therefore, this example allows us to quantify the effect of our approximation while still controlling the score estimation error (through $\epsilon$). The following results evaluate the robustness of the algorithms to increasing error rates $\epsilon$ and number of conditioning observations $n$.
@@
Table 1 displays the total running time and sW [^3] The sW is computed with $10^{4}$ slices on $10^{3}$ samples. To account for finite-sample effects, we considered a normalized version of sW by removing the expected sW over different sample sets from the true distribution. for each algorithm on the Gaussian example (Table 3 in Appendix G shows similar results for GMM). We can see that for the same number of time steps, our algorithm yields smaller sW than the Langevin sampler while accounting for $L=5$ times less neural network evaluation, thus $5$ times faster. We show in Appendix G the differences in speed for each tested setting. Based on this table, we consider “equivalent time settings”, namely we run our algorithm with $400$ and $1000$ steps for JAC and GAUSS respectively and LANGEVIN for $400$ steps.
@@
Figure 2: Sliced Wasserstein (sW) distance as a function of the number of observations $n$ for the Gaussian and GMM toy examples, for each algorithm (GAUSS, JAC and LANGEVIN) and with different levels of $\epsilon$. Mean and std over 5 different seeds.